The Fan Cost Index​ (FCI) is a measure of how much it costs a family of four to attend a major league baseball game.

The frequency distribution and the histogram for the Fan Cost Index are given to the right.

FCI​ (Dollars)

Frequency

​100-149.99

44

​150-199.99

1313

​200-249.99

99

​250-299.99

11

​300-349.99

11

​350-399.99

00

​400-449.99

22

10020030040050002468101214Fan Cost IndexFrequency

A bar graph has a horizontal axis labeled Fan Cost Index from 50 to 500 in increments of 50 and a vertical axis labeled Frequency from 0 to 14 in increments of 2. There are vertical bars each with width 50 and all over the horizontal axis. The cost and frequency follow, with the cost first: 100-150, 4; 150-200, 13; 200-250, 9, 250-300, 1; 300-350, 1; 350-400, 0; and 400-450, 2.

​(a) Determine the shape of the distribution.

A.

Right Skewed Right Skewed

B.

Normal

C.

Left SkewedLeft Skewed

D.

Uniform

​(b) Determine if the mean or the median would be a better measure of center for this distrubution. If the mean is​ better, calculate the value. If the median is​better, give the midpoint of the interval that contains it.

Select the correct choice below and fill in the answer box to complete your choice.

A.

The mean is a better measure of center for this distribution.

The value of the mean is

​$nothing.

​(Round to the nearest integer as​ needed.)

B.

The median is a better measure of center for this distribution.

The midpoint of the interval that contains the median is

​$nothing.

​(Round to the nearest integer as​ needed.)

Click to select and enter your answer(s).
Answer #1

125 4218642 0 A3uanbau

a) as we can see graph ,data is right skewed

so, option a) is answer

b)

since graph is skewd ,so median will be better representative for measure of center for this distrubution

Total number of frequency=30

so, median = (30+1)/2 = 15.5th term

FCI​ (Dollars) mid point Frequency cumulative
​100-149.99 125 4 4
​150-199.99 175 13 17
​200-249.99 225 9 26
​250-299.99 275 1 27
​300-349.99 325 1 28
​350-399.99 375 0 28
​400-449.99 425 2 30

So median lies in the interval 150 - 199.9.

mid point = (150+200)/2 = 175

answer- option b

The median is a better measure of center for this distribution.

The midpoint of the interval that contains the median is 175